OpenStudy (anonymous):

P has degree 3 and zeros 4 and i.

OpenStudy (anonymous):

(X+5)/2=1 x=-3 (Y+5)/2=4 Y=3 Point is (-3,3)

OpenStudy (anonymous):

not solving for a point, solving for equation P(x)

Directrix (directrix):

Zeros are 4, i, and - i (x - 4) (x - i) ( x - (-i)) = P(x) (x - 4) ( x - i) (x + i) = P(x) x^3 - 4 x^2 + x - 4 = P(x)

OpenStudy (anonymous):

Q has degree 3, and zeros 0 and i

OpenStudy (anonymous):

x(x-i)(x+i)= x^3+x?